The displacement time graph for two particles A and B are straight lines inclined at angles of 30o and 60o with the time axis . The velocities of V: V is ?

Grade:11

3 Answers

Arun
5 years ago
Velocity = Slope of the line formed in displacement v/s time graph = Tanθ
Va : Vb = Tan30 : Tan60 = (1/√3) :√3= 1 : 3
Ratio of velocities of A:B is 1 : 3
 
Regards
Arun (askIITians forum expert)
Khimraj
5 years ago
velocity is slope of displacement time graph.
SO
Va/Vb = tan30/tan60 = 1/3.
Hope it clears.
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
Velocity is slope of displacement time graph.
So,
Va/Vb = tan30/tan60 = 1/3.
or, Va:Vb = 1:3
 
Thanks and regards,
Kushagra